OpenStudy (anonymous):

What is the value of expression? 5^−8 • 5^−4 A. 5^32 B. 5^4 C. 5^−4 D. 5^−12

OpenStudy (anonymous):

A

geerky42 (geerky42):

\[\Large a^b\cdot a^c = a^{b+c}\]

OpenStudy (fibonaccichick666):

So, what if you had \((5^2)^3\) what would that become? how about \(5^2*5^3\), what process do you use?

OpenStudy (anonymous):

the answer is A

geerky42 (geerky42):

Explain why @orangejuicegood

OpenStudy (anonymous):

because it is

geerky42 (geerky42):

no explanation, so I assume it is wrong.

OpenStudy (fibonaccichick666):

Ignore the troll, @themanplan , @geerky42

OpenStudy (anonymous):

Wow okay, rude.

geerky42 (geerky42):

it's like saying 1+1=3 just because it is that way.

OpenStudy (anonymous):

it equals 2

OpenStudy (fibonaccichick666):

that would be the definition of trolling^^

OpenStudy (anonymous):

how do you expect to help people with math if you cant do 1+1

geerky42 (geerky42):

you missed my point...

geerky42 (geerky42):

i really hope you are trolling...

OpenStudy (anonymous):

do know what that means...

OpenStudy (anonymous):

sounds offensive

OpenStudy (anonymous):

so the answer not a

OpenStudy (fibonaccichick666):

no, but if you answer geerky or myself, we will help you

OpenStudy (anonymous):

no it is a I 100% guarantee it

geerky42 (geerky42):

then provide your work

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

np

OpenStudy (anonymous):

was it correct ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

good :)

geerky42 (geerky42):

And..... troller won for this round

OpenStudy (anonymous):

your just bad at math, i can tutor you if you like @geerky42

geerky42 (geerky42):

it's not me being bad at math, it's you being bad at reading comprehension

geerky42 (geerky42):

like I said, you missed my point...

OpenStudy (anonymous):

okay